40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 23 20:58:38 2014, ericq, Update, CDS, Bootfest 2014! 
    Reply  Tue Jun 24 01:07:56 2014, Jenne, Update, LSC, Bootfest 2014! MMT1_face_23June2014_2315pm.png
       Reply  Tue Jun 24 18:35:53 2014, Jenne, Update, LSC, Still no beam going to IFO 
Message ID: 10090     Entry time: Tue Jun 24 01:07:56 2014     In reply to: 10088     Reply to this: 10094
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Bootfest 2014! 

Still no real luck getting the beam back aligned to the IFO.

Koji and I tried a few minutes of wiggling the input pointing tip tilts (TT1 and TT2) around, and then tried doing some thinking.

We note that the beam propagates (modulo a few pickoffs):

IMC -> Faraday -> TT1 -> MMT1 -> MMT2 -> TT2 -> PRM.

Since moving TT1 to the rails does make beam reflections in the BS chamber move (as seen by movement of the general illumination on the PRM face camera), I posit that the beam is getting through the Faraday.  It is certainly getting at least mostly through the Faraday, although since the MC locked so easily, I assume that we didn't have too much movement after the ~2pm Alaskan earthquake & aftershocks, so we're at pretty much the same alignment as usual, in terms of beam pointing coming from the IMC.

The plan is then to see the position of the beam on MMT1, and steer using TT1 to get the beam to roughly the center.  Then, see the beam propagate to MMT2 (if possible) and TT2 (if possible).    From here, we should be able to see the spot on PRM.  We should be able to use TT2 to tweak things up, and get the beam back to about the right place on POP, or REFL, or somewhere farther along.  Hopefully at this point, we'd see some flashes in the Yarm. 

Using a spare Watek camera, I was able to capture a shot of the face of MMT1.  This is when the TTs were restored to their values that were saved last Monday.  I checked, and this is also roughly the center of the actuation range of TT1, for both pitch and yaw.

MMT1_face_23June2014_2315pm.png

I am not able to see the face of MMT2, or TT2.  If I leave TT1, and move TT2, I am not able to see any movement of any beam or reflections seen in the PRM face camera.

Koji and I are checking the MC spot positions, but it may be time to leave this for the morning crew.

EDIT:  The MC spots were actually pretty bad, and the WFS were working really hard.  Koji realigned the MC suspensions, and now the MC spots are slightly better, although quite similar, to what Manasa measured last week.  The restored TT values still don't give us any flashes in the arms.

ELOG V3.1.3-